OpenStudy (anonymous):

The water lily Growerosa Veryfasterosa doubles its size every day. From the time it was planted until it completely covered the pond took 26 days.

OpenStudy (anonymous):

and the question is?

OpenStudy (anonymous):

How many days did it take to cover half the pond?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

since it doubles every day, if it was full on day 26 it was half full on day 25, one quarter full on day 24, one eighth full on day 23 etc

OpenStudy (anonymous):

what is the last digit of 3^100 and is there an easy way to find it?

OpenStudy (anonymous):

depend on what number theory you know. but you can do it by looking at a pattern \(3^0=1,3^2=9,3^3=27,3^4=81,3^5=243\) etc

OpenStudy (anonymous):

so last digit is either 1,3,9,7 and it repeats in that order

OpenStudy (anonymous):

if the exponent is 4, 8, 12, 16, ... i.e. a multiple of 4, you get the last digit is 1. since 4 goes in to 100 evenly, the last digit will also be 1

OpenStudy (anonymous):

If the rabbit population of an island is given by R=1500x( 0.817)^t where t is the time in weeks after a virus was introduced to eradicate them, can you find: a. the intial population b. the population after 31/2 weeks